# Matching seismic masses for RR Lyrae-type and oscillating red horizontal-branch stars in M4

### v1

@@ 4.2 Seismic masses for red giants and the average rHB mass
Pressure modes in solar-like oscillators – such as red giants – are characterized by two global seismic parameters; the large frequency separation, , and the frequency of the maximum acoustic power, . These quantities are correlated to stellar properties, which are used to derive seismic mass scaling relations (Ulrich, 1986; Brown et al., 1991; Kjeldsen & Bedding, 1995). For time-series photometry with short baselines (e.g. the K2 mission), it has been shown that accurate asteroseismic masses can be measured independently of a estimate (e.g. Howell et al., 2022, 2024). This method uses the following scaling relation:
@@
We use the red giant sample and their corresponding K2 light curves from Howell et al. (2022) to measure seismic masses using this relation above. They detected solar-like oscillations in 75 red giants across three phases of evolution (RGB, rHB and early AGB), and used the pySYD pipeline (Chontos et al., 2022) to measure the asteroseismic parameters. We adopt the Howell et al. (2022) estimates for and luminosity, although not their determinations. Here we report updated measurements of and the corresponding uncertainty for each star using a new asteroseismic pipeline, pyMON (Howell et al., in prep).
@@
The pyMON pipeline is an adaptation of pySYD (see Chontos et al. 2022 for more details), where the main difference is that is not measured. In both pipelines, the power spectrum is smoothed using an estimate for . pyMON substitutes this measurement with an estimate from a calibrated - scaling relation (Stello et al., 2009; Howell et al., 2022). This is useful for low signal-to-noise data (e.g. our M4 photometry), because it is difficult to measure an accurate . A comparison between the pySYD and pyMON methods shows that the measurement of remains consistent, however there is a reduction by a factor of in the uncertainty, due to the reduced uncertainty in . Consequently, this results in a decrease in the seismic mass uncertainties by a factor of .
We provide our individual seismic masses in the bottom panel of Figure 6. Note that we have removed the stars identified as mass outliers in Howell et al. (2022), and also limited the RGB sample to below the luminosity bump (hereby known as lower RGB or LRGB). This ensured there was no bias to lower masses due to any mass loss that might have occurred. In the top panel of Figure 6, we represent the distribution of our masses as kernel density estimation (KDE) functions for each evolutionary phase. A KDE is calculated by modelling each data point as a Gaussian, adopting the mass as the mean and the uncertainty as the standard deviation. The final KDE is the summation of the Gaussian components, where the area has been normalised to a value of one.
@@
To estimate a new average mass for the HB sample, we determine the peak value of the corresponding KDE. This represents the mode mass for this sample, and was measured to be The uncertainty is calculated as the standard error on the mean. Our average HB mass is consistent with the value found previously in Howell et al. (2022). This was similarly found for the RGB and early AGB samples, where we calculated average masses of and respectively444Note Howell et al. (2022) used the mean mass as their average mass scale. In this paper, we adopt the same method as Howell et al. (2024) of measuring the mode mass of the KDE distributions to determine the average mass, leading to small discrepancies. However, both average mass measures are consistent within uncertainties.. Individual and average masses relative to the calculated luminosities are plotted in Fig. 7.
The integrated mass loss rates between evolutionary phases are also consistent with Howell et al. (2022). The mass loss from the lower RGB to the HB decreases slightly to , whereas we find a mass loss from from the HB to the early AGB.
We plot the distributions of the masses along the HRD in Fig. 8. We also include a MIST555https://waps.cfa.harvard.edu/MIST (MESA Isochrones and Stellar Tracks; Choi et al. 2016) isochrone for comparison. This isochrone has an age of 12.2 Gyr, an average of various literature sources for M4 (Cabrera-Ziri & Conroy, 2022). Since MIST does not offer -enhanced isochrones, we use a value of [Fe/H] = –0.9 to emulate the higher [/Fe] abundances, based on the scaling described by Joyce et al. (2023). The mass loss prescriptions used by MIST for low-mass stars are and for the RGB and AGB phases, following the (Reimers, 1975) and Bloecker (1995) schemes, respectively. Colors represent masses: it is clear that the conservative mass loss prescription used in MIST’s pre-computed isochrone database does not capture the true amount of mass loss in the cluster. This is most evident in the later evolutionary stages: while the isochrone has masses around 0.82–0.83 in the RGB, this only drops to 0.79 by the time of the HB and early AGB, which is significantly higher than the observed values. A consequence of the overly conservative mass loss prescription is that the isochrone barely reaches the red edge of the RHB region and therefore cannot reproduce the structure of the HB at all. We note that mass loss parameters cannot be adjusted in MIST through the web interface, either.
@@ 4.3 Comparison of the rHB and RRc masses
For M4, we found a combined HB mass of M⊙. If we split this sample into their constituent RR Lyrae and rHB parts, we find average masses of M⊙ and M⊙, respectively. A small non-zero difference is in agreement with the evolutionary predictions that stellar (envelope) masses along the HB decrease towards the blue (Salaris & Cassisi, 2006). Therefore, we can conclude that our two seismic mass estimates, the scaling relations for rHB stars, and the linear RRc model grid of Netzel & Smolec (2022), are in agreement. Thus, we may conclude that fitting the modes offers a reliable way to estimate RR Lyrae star masses, even if it is only available for overtone stars.
4.4 Comparison with mass relations
@@
With these limitations in mind, we calculated the RR Lyrae masses using these relations and compared them to our seismic results in two ways. First we used the classical constraints, and values, computed from the Gaia photometry, which also served as initial values for our seismic model fits. These results are plotted in the left panel of Fig. 9. As the figure shows, the values are generally in the right range, between 0.55 and 0.8 , however, differences from the seismic values can reach 0.1 , and uncertainty ranges can be even wider than that, ranging from to .
@@
Gratton et al. (2010) determined a different mass relation based on the (or ) color indices and [Fe/H] indices of HB stars. We calculated the masses of both the RRc and the rHB stars using our dereddened colors. The results are displayed in Fig. 10. This relation results in a larger mass difference (0.037 ) between the two groups than what we found from the seismic data. We suspect that this difference is coming from the fact that the relation is based on zero-age HB stellar models, whereas the stars in M4 are not necessarily on the ZAHB, but rather at evolutionary stages dictated by their common ages and thus common masses.
@@ 5 Conclusions
Globular clusters, such as M4, provide an opportunity to probe the late stages of low-mass stellar evolution. Building on the unique capabilities of the Kepler space telescope, we calculated new, more accurate masses for the oscillating giant stars using asteroseismic scaling relations and values determined with the new pyMON code (Howell et al., in prep.). We then analysed the overtone RR Lyrae stars in the cluster and determined their masses using peak-bagging and fitting the frequencies with the seismic model grid of Netzel & Smolec (2022). With these, we were able to relate RR Lyrae seismic masses to seismic masses of red HB stars, and found that they match each other closely. Furthermore, we estimate a He content of for the cluster from our models.
While our result is not yet a direct measurement of RR Lyrae masses, as it relies on the assumption that RR Lyrae and rHB stars in the cluster have similar masses, we were able to test model-dependent seismic (or pulsational) masses against independent mass estimates for the first time. We find that the average RRc and rHB masses agree within the errors, with the RRc stars having slightly lower masses, as expected from the stucture of the HB. We compared the modeled physical parameters of the RRc stars to various relations between M, L, and the pulsation periods, and found that while the generally agree, systematical differences do exist. We also show, however, that the usability of these relations is rather limited, as L and values have to be determined very accurately, otherwise the calculated masses will be very uncertain.
This study not only tests RR Lyrae masses for the first time, but it also lends further support to the hypothesis of the modes being high-order modes, as proposed by Dziembowski (2016). Therefore, it is now possible to estimate masses for field RRc and RRd stars as well, where enough modes can be detected from high-precision photometry (Netzel & Smolec, 2022; Netzel et al., 2023). Furthermore, RRc stars now offer a way to measure masses on the HB in globular clusters, even where rHB stars are too faint for asteroseismology. RR Lyrae stars thus open up a new way to study mass loss in more clusters, either with Kepler, or with future instruments, such as the Roman Space Telescope (Howell et al., 2024; Molnár et al., 2023a) or the HAYDN telescope project, which is designed specifically to observe dense stellar fields (Miglio et al., 2021).
@@ Acknowledgements.
This research was supported by the ‘SeismoLab’ KKP-137523 Élvonal grant of the Hungarian Research, Development and Innovation Office (NKFIH). Cs.K. was supported by the ÚNKP-23-3, New National Excellence Program of the Ministry of Culture and Innovation from the source of the National Research, Development and Innovation Fund. This research has received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 947660). H.N. is funded by the Swiss National Science Foundation (award PCEFP2_194638). M.J. gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commission’s Widening Fellowship. This project has received funding from the European Union’s Horizon 2020 research and innovation programme. This research made use of NASA’s Astrophysics Data System Bibliographic Services, as well as of the SIMBAD database operated at CDS, Strasbourg, France.References
@@
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
@@
- Mochejska et al. (2002) Mochejska, B. J., Kaluzny, J., Thompson, I., & Pych, W. 2002, AJ, 124, 1486
- Molnár et al. (2023a) Molnár, L., Kalup, C., & Joyce, M. 2023a, arXiv e-prints, arXiv:2306.12459
@@
- Momany et al. (2002) Momany, Y., Piotto, G., Recio-Blanco, A., et al. 2002, ApJ, 576, L65
- Mucciarelli & Bellazzini (2020) Mucciarelli, A. & Bellazzini, M. 2020, Research Notes of the American Astronomical Society, 4, 52
@@
- Netzel et al. (2024) Netzel, H., Varga, V., Szabo, R., Smolec, R., & Plachy, E. 2024, arXiv e-prints, arXiv:2404.14373

### v3

@@ 4.2 Seismic masses for red giants
Mixed mode oscillations in red giants are characterized by two global seismic parameters; the large frequency separation, , and the frequency of the maximum acoustic power, . These quantities are correlated to stellar properties, which are used to derive seismic mass scaling relations (Ulrich, 1986; Brown et al., 1991; Kjeldsen & Bedding, 1995). For faint globular cluster stars, the short observing baseline of the K2 mission results in lower frequency resolution and signal-to-noise, and as such it can be difficult to estimate . However, accurate asteroseismic masses can be measured independently of a estimate (e.g. Howell et al., 2024, 2025; Reyes et al., 2025), using the following scaling relation:
@@
We use the oscillating red giant sample and the corresponding K2 light curves from Howell et al. (2022) to measure seismic masses using Eq. 4. They detected solar-like oscillations in 75 red giants across three phases of evolution (RGB, rHB and early AGB), and used the pySYD pipeline (Chontos et al., 2022) to measure the asteroseismic parameters. We adopt the Howell et al. (2022) estimates for and luminosity, although not their measurements. Here we report updated measurements of and the corresponding uncertainty for each star using a new asteroseismic pipeline, pyMON (Howell et al., 2025).
@@
The pyMON pipeline is an adaptation of pySYD (see Chontos et al. 2022 for more details), where the main difference is that is not measured. In both pipelines, the power spectrum is smoothed using an estimate for , however pyMON uses a calibrated - scaling relation (Stello et al., 2009) to estimate this quantity. This is useful for low signal-to-noise data – such as the M4 red giant sample in Howell et al. (2022)– where measurements are are highly uncertain. A comparison between the pySYD results in Howell et al. (2022) and our pyMON results show that the measurement of remains consistent, however there is a reduction by a factor of in the uncertainty (refer to Howell et al. 2025 for further discussion on this uncertainty reduction). Consequently, this results in a decrease in the seismic mass uncertainties by a factor of .
Scaling relations may also include correction factors (, ) to account for structural differences between the Sun and the target stars. In this case we are only interested in the value of , for which various prescriptions have been given in the literature (see, e.g., Sharma et al., 2016; Hekker, 2020; Pinsonneault et al., 2025), but these have been limited to higher metallicities. More recently, Lindsay et al. (2025) found that at the metallicity of M4, , which could indicate an overestimation of about 5% of RGB masses. However, that only applies to RGB stars, and correction factors have not been estimated for low-metallicity core-He-burning or AGB stars yet.
We note, however, that correction factors are predominantly calibrated through (or involving) stellar radii, and Ash et al. (2025) showed that radius-based values could potentially result in overcorrecting the masses. They also found that while not using any corrections can result in small systematic errors in the seismic masses, the differences stay largely constant along the RGB (and larger, -dependent deviations only occur when is also used). Thus, if does not depend on , the mass differences between different stars will not be affected.
Furthermore, clusters offer the possibility to test the masses against stellar models and isochrones. Howell et al. (2022) found that the RGB masses agree with the masses of the fitted isochrone without the need for correction factors. We also show the agreement between our RGB masses and stellar models in Sect. 4.4. We also note that decreasing our AGB masses by 5% would be in tension with the white dwarf masses in the cluster (see Sect. 4.3). Based on these arguments, and also to stay consistent with analyses of other globular clusters which used (Howell et al., 2022, 2024, 2025), we decided to keep at unity.
4.3 Average masses of the subpopulations
Most Galactic globular clusters contain multiple populations of stars, usually signaling two consecutive waves of star formation (Milone & Marino, 2022). M4 is no exception to this: the possible signs of both populations in the RGB data and their mass difference was already explored by Howell et al. (2022). However, since the RGB population showed only weak signs of bimodality at the mass, disentangling the RGB population will require further chemical abundance measurements. Since here we are more concerned about the later stages of stellar evolution, we treat the RGB sample as a single population. Using spectroscopy, MacLean et al. (2018) reported that all AGB stars in M4 belong to the first generation of stars, with second-population stars possibly skipping the AGB altogether. This suggests that the EAGB stars in Howell et al. (2022) only belong to the first generation. Most importantly, Marino et al. (2011) found that both the rHB and RR Lyrae stars of the cluster are first-population stars, and the distinct blue arm of the HB contains the second population. This makes it possible to compare their masses to each other and to the EAGB stars, too.
@@ 4.3 Average masses of the subpopulations
In Figure 6, we show the distribution of our masses as kernel density estimation (KDE) functions (top panel) and as a scatter plot (bottom panel) for each evolutionary phase. We removed the stars identified as mass outliers in Howell et al. (2022) and our study (V61). Additionally, we limited the RGB sample to below the luminosity bump (hereby known as lower RGB or LRGB) to ensure there was no bias to lower masses due to any mass loss that might have occurred. To estimate a new average mass for the rHB sample, we determine the peak value of the corresponding KDE. This represents the mode mass for this sample, and was measured to be , where the uncertainty is calculated as the standard error on the mean. Our average rHB mass is consistent with the value found previously in Howell et al. (2022). This was similarly found for the RGB and early AGB samples, where we calculated average masses of and respectively444We highlight that Howell et al. (2022) used the mean mass as their average mass scale. In this paper, we adopt the same method as Howell et al. (2024) of measuring the mode mass of the KDE distributions to determine the average mass. Both average mass measures are consistent within uncertainties.. Individual and average masses relative to the calculated luminosities are plotted in Fig. 7, and listed in Table 3.
@@
The integrated mass loss between evolutionary phases are also consistent with Howell et al. (2022). The mass loss from the lower RGB to the HB decreases slightly to , whereas we find a mass loss from from the HB to the early AGB.
A further mass constraint for the cluster was determined by Kalirai et al. (2009), who inferred masses of 0.50–0.55 M⊙ for six white dwarfs. As discussed by Howell et al. (2022), this is very close to the AGB seismic masses, indicating very low mass loss on the AGB, which can be explained by the stars skipping most or all of the thermally pulsing stage of the AGB. We note that discrepancies between AGB evolutionary models and elemental abundances in the AGB stars have also been reported MacLean et al. (2018). A KDE-based fitting of the WD data, consistent with our analysis of the other populations, resulted in an average mass of M⊙, which is lower than the average published by Kalirai et al. (2009). While this is still quite close to the EAGB mass, the difference is 0.03 M⊙, suggesting a small amount of AGB mass loss. We note, however, that the AGB mass distribution is based on only four stars, which makes the AGB average mass rather uncertain.
4.4 The seismic HRD of M4
We plot the distributions of the masses along the HRD in Fig. 8. We also include a MIST555https://waps.cfa.harvard.edu/MIST (MESA Isochrones and Stellar Tracks; Choi et al. 2016) isochrone for comparison. This isochrone has an age of 12.2 Gyr, an average of various literature sources for M4 (Cabrera-Ziri & Conroy, 2022). Since MIST does not offer -enhanced isochrones, we use a value of [Fe/H] = –0.9 to emulate the higher [/Fe] abundances, based on the scaling described by Joyce et al. (2023). The mass loss prescriptions used by MIST for low-mass stars are and for the RGB and AGB phases, following the (Reimers, 1975) and Bloecker (1995) schemes, respectively. Colors represent masses: it is clear that the conservative mass loss prescription used in MIST’s pre-computed isochrone database does not capture the true amount of mass loss in the cluster. This is most evident in the later evolutionary stages.
The isochrone has masses around 0.82–0.83 in the RGB, which agrees with our inferred LRGB masses to within 3%, even without further optimization of the isochrone fit. We note that this again indicates that no correction is necessary for the scaling relation, since literature values would decrease the seismic masses. However, the isochrone mass only drops to 0.79 by the time of the HB and early AGB, which is significantly higher than the observed values. A consequence of the overly conservative mass loss prescription is that the isochrone barely reaches the red edge of the rHB region and therefore cannot reproduce the structure of the HB at all. We note that mass loss parameters cannot be adjusted in MIST through the web interface, either.
In order to capture the structure of the HB, we also included two zero-age horizontal branch (ZAHB) tracks in Fig. 8 from the Bag of Stellar Tracks and Isochrones (BaSTI) model library (Pietrinferni et al., 2021). Both tracks use alpha-element enhanced models ([/Fe] = 0.4). The two pre-computed tracks closest to the [Fe/H] index of M4 have indices of –0.9 and –1.2, respectively. We also tested alpha-enhanced BaSTI isochrones at the same age as the MIST one in the plot, but they only differed significantly from MIST at the main sequence turnoff and were nearly indistinguishable along the giant branch.
@@ 4.5 Comparison of the rHB and RRc masses
For M4, we found a combined HB mass of M⊙. If we split this sample into their constituent RR Lyrae and rHB parts, we find average masses of M⊙ and M⊙, respectively. The small difference of M⊙ is in agreement with the evolutionary predictions that stellar (envelope) masses along the HB decrease towards the blue (Salaris & Cassisi, 2006). Therefore, we can conclude that our two seismic mass estimates, the scaling relations for rHB stars, and the linear RRc model method of Netzel & Smolec (2022), are in agreement. Thus, we may conclude that fitting the modes offers a reliable way to estimate RR Lyrae star masses, even if it is only available for overtone stars.
4.6 Comparison with mass relations and models
@@
With these limitations in mind, we calculated the RR Lyrae masses using these relations and compared them to our seismic results in two ways. First we used the classical constraints, and values, computed from the Johnson photometry, which also served as initial values for our seismic model fits. These results are plotted in the left panel of Fig. 9. As the figure shows, the values are generally in the right range, between 0.45 and 1.0 , however, differences from the seismic values can reach 0.1–0.3 .
@@
Gratton et al. (2010) determined a different mass relation based on the (or ) color indices and [Fe/H] indices of HB stars. We calculated the masses of both the RRc and the rHB stars using our dereddened colors. The results are displayed in Fig. 10. This relation results in a larger mass difference (0.037 ) between the two groups than what we found from the seismic data. This difference could be related to he fact that the relation is based on zero-age HB stellar models, whereas the stars in M4 are not necessarily on the zero-age horizontal branch (ZAHB), but rather at evolutionary stages dictated by their common ages and thus common masses.
Finally, we compared our results directly to the BaSTI ZAHB models. We plot three model sequences in Fig. 11 on the – mass plane. All three models are -enhanced ([/Fe] = +0.4), with differing [Fe/H] indices (-1.3, -1.2, -0.9) but similar He contents (Y = 0.259, 0.250, 0.252). While the inferred mass and values do not line up well with a single track, three out of four RRc stars within from the measured [Fe/H] = value of the cluster, with RRc stars aligning with slightly more metal-poor (more massive), and RHB stars with slighly metal-rich (lighter) sequences. This result agrees with what we observe in Fig. 10. However, whether the ZAHB models overestimate the mass gradient along the HB or our mass measurements contain systematic offsets cannot be decided from these results alone.
@@ 5 Conclusions
Globular clusters, such as M4, provide an opportunity to probe the late stages of low-mass stellar evolution. Building on the unique capabilities of the Kepler space telescope, we calculated new, more accurate masses for the oscillating giant stars using asteroseismic scaling relations and values determined with the new pyMON code (Howell et al., in prep.). We then analysed the overtone RR Lyrae stars in the cluster and determined their masses via peak-bagging and fitting the frequencies with seismic model calculations Netzel & Smolec (2022); Netzel et al. (2023). With these, we were able to relate RR Lyrae seismic masses to seismic masses of red HB stars, and found that they match each other closely, with the RR Lyrae models being lighter, as expected from stellar evolution. We found the average rHB and RRc masses to be M⊙ and M⊙, respectively.
Furthermore, we estimate a low He enrichment in the cluster, in agreement with the findings of Villanova et al. (2012), based on the RRc model results. However, the models leave the chemical compositions ambiguous for multiple targets, which makes this result uncertain.
While our result is not yet a direct measurement of RR Lyrae masses, as it relies on the assumption that RR Lyrae and rHB stars in the cluster have very similar masses, we were able to test model-dependent seismic (or pulsational) masses against independent mass estimates for the first time. We find that while the RRc stars have slightly lower masses, as expected from the stucture of the HB, the difference of M⊙ is less than what stellar evolutionary models predict. Whether the source of this discrepancy comes from the pulsational or the evolutionary models remains to be seen: but even if considering such systematic differences, we estimate that our method to determine RR Lyrae masses is accurate to M⊙.
We also compared the modeled physical parameters of the RRc stars to various relations between M, L, and the pulsation periods, and found that while these generally agree, systematical differences do exist. We also show, however, that the usability of these relations is rather limited, as L and values have to be determined very accurately, otherwise the calculated masses will be very uncertain.
This study not only tests RR Lyrae masses for the first time, but it also lends further support to the hypothesis of the modes being high-order modes, as proposed by Dziembowski (2016). Therefore, it is now possible to estimate masses (and values) for field RRc and RRd stars to within a few hundredth M⊙ accuracy, if enough modes can be detected from high-precision photometry (Netzel & Smolec, 2022; Netzel et al., 2023). Furthermore, RRc stars now offer a way to measure masses on the HB in globular clusters, even where rHB stars are too faint for asteroseismology. RR Lyrae stars thus open up a new way to study mass loss in more clusters, either with Kepler (Kalup et al., in prep.), or with future instruments, such as the Roman Space Telescope (Molnár et al., 2023a) or the HAYDN telescope project, which is designed specifically to observe dense stellar fields (Miglio et al., 2021).
@@ Acknowledgements.
This research was supported by the ‘SeismoLab’ KKP-137523 Élvonal grant and the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH), and by the LP2025-14/2025 Lendület grant of the Hungarian Academy of Sciences. Cs.K. was supported by the ÚNKP-23-3, New National Excellence Program of the Ministry of Culture and Innovation from the source of the National Research, Development and Innovation Fund. This research has received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 947660). H.N. is funded by the Swiss National Science Foundation (award PCEFP2_194638) and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 951549 - UniverScale). M.J. gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commission’s Widening Fellowship. This project has received funding from the European Union’s Horizon 2020 research and innovation programme. This research made use of NASA’s Astrophysics Data System Bibliographic Services, as well as of the SIMBAD database operated at CDS, Strasbourg, France.References
- Abdollahi et al. (2025) Abdollahi, H., Molnár, L., & Varga, V. 2025, A&A, 695, L14
@@
- Alonso et al. (1999) Alonso, A., Arribas, S., & Martínez-Roger, C. 1999, A&AS, 140, 261
- Alonso-Garc´ıa et al. (2012) Alonso-García, J., Mateo, M., Sen, B., et al. 2012, AJ, 143, 70
@@
- González Hernández & Bonifacio (2009) González Hernández, J. I. & Bonifacio, P. 2009, A&A, 497, 497
@@
- Hekker (2020) Hekker, S. 2020, Frontiers in Astronomy and Space Sciences, 7, 3
@@
- Howell et al. (2025) Howell, M., Campbell, S. W., Kalup, C., Stello, D., & De Silva, G. M. 2025, MNRAS, 536, 1389
@@
- Johnson et al. (2014) Johnson, C. I., Rich, R. M., Kobayashi, C., Kunder, A., & Koch, A. 2014, AJ, 148, 67
@@
- Lindsay et al. (2025) Lindsay, C. J., Hon, M., Ong, J. M. J., et al. 2025, arXiv e-prints, arXiv:2507.01091
- Liška et al. (2016a) Liška, J., Skarka, M., Mikulášek, Z., Zejda, M., & Chrastina, M. 2016a, A&A, 589, A94
- Liška et al. (2016b) Liška, J., Skarka, M., Zejda, M., Mikulášek, Z., & de Villiers, S. N. 2016b, MNRAS, 459, 4360
@@
- Milone & Marino (2022) Milone, A. P. & Marino, A. F. 2022, Universe, 8, 359
- Molnár et al. (2023a) Molnár, L., Kalup, C., & Joyce, M. 2023a, NASA Roman White Papers, arXiv:2306.12459
@@
- Netzel et al. (2024) Netzel, H., Varga, V., Szabó, R., Smolec, R., & Plachy, E. 2024, A&A, 692, A133
@@
- Pancino et al. (2024) Pancino, E., Zocchi, A., Rainer, M., et al. 2024, A&A, 686, A283
- Pietrinferni et al. (2021) Pietrinferni, A., Hidalgo, S., Cassisi, S., et al. 2021, ApJ, 908, 102
@@
- Sharma et al. (2016) Sharma, S., Stello, D., Bland-Hawthorn, J., Huber, D., & Bedding, T. R. 2016, ApJ, 822, 15
- Skrutskie et al. (2006) Skrutskie, M. F., Cutri, R. M., Stiening, R., et al. 2006, AJ, 131, 1163
@@
- Stetson et al. (2019) Stetson, P. B., Pancino, E., Zocchi, A., Sanna, N., & Monelli, M. 2019, MNRAS, 485, 3042
- Sylla et al. (2024) Sylla, S., Kolenberg, K., Klotz, A., et al. 2024, A&A, 691, A108
@@
- Tailo et al. (2022) Tailo, M., Corsaro, E., Miglio, A., et al. 2022, A&A, 662, L7
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Heinze, A. N., et al. 2018, PASP, 130, 064505
@@
- Valcarce et al. (2014) Valcarce, A. A. R., Catelan, M., Alonso-García, J., Cortés, C., & De Medeiros, J. R. 2014, ApJ, 782, 85